# Operation of silicon photomultipliers in a dilution refrigerator down to 9.4 mK towards a cryogenic cosmic ray muon veto system

### v1

@@ 3.3 Direct crosstalk
Optical crosstalk is a source of correlated noise in a SiPM that occurs due to the emission of a small number of photons ($\sim$10-5 photons/electron) during an avalanche [mirzoyan2009light], which can trigger avalanches in the surrounding microcells. Optical crosstalk can be split into two categories:
@@
DiCT manifests as a superposition with the 1 PE (single photon) pulse. For example, if a primary avalanche triggers two secondary crosstalk avalanches, the resulting output pulse will be a 3 PE pulse. DiCT can therefore be a limiting factor for applications requiring accurate photon-counting, since it artificially increases the number of detected photons and degrades the signal-to-noise ratio. The DiCT probability is expected to increase with $\Delta V$, since the probability of an avalanche increases with $\Delta V$. In addition, the number of emitted photons is linearly proportional to $\Delta V$, which also contributes to an increase in the DiCT probability. There appears to be a weak dependence on the DiCT probability with temperature for the NUV-HD-cryo SiPM technology, as reported in [NUV-HD-cryo].
To characterise the DiCT of the SiPM, we calculate the mean number of additional prompt avalanches per primary, ($\bar{N}_{\mathrm{APA}}$), defined as:
$\bar{N}_{\mathrm{APA}}=\frac{1}{N}\sum_{i=1}^{N}{\frac{Q_{i}}{\bar{Q}_{\mathrm{1PE}}}}-1,\qquad$ (3.2)
where $Q_{i}$ is the charge of each identified pulse, $\bar{Q}_{\mathrm{1PE}}$ is the average 1 PE charge, and $N$ is the number of pulses analysed. In practise, we compute $\bar{N}_{\mathrm{APA}}$ by normalising the amplitude distribution shown in Figure 5 by the average 1 PE amplitude, $\bar{A}_{\mathrm{1PE}}$, to align the peaks on the $x$-axis with the number of PE they correspond to. We then take the mean of the normalised histogram and subtract 1 to determine $\bar{N}_{\mathrm{APA}}$, shown in Figure 9 for an operating voltage of $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 4.8 V).
We choose the amplitude distribution instead of the charge distribution for this analysis due to the fact that for increasing $\Delta V$, we start to observe an additional “shoulder” feature forming in the tails of the peaks in the charge distribution. This is an artifact of the charge integration window used in this analysis: if correlated delayed avalanches (CDAs) following an avalanche occur within the charge integration window of the pulse of interest, $t_{p}-100$ ns : $t_{p}+250$ ns], part or all of the CDA pulse(s) will be included in the charge integral of the pulse of interest, skewing the charge distribution. Figure 9 (right) shows the pulse charge distribution of the NUV-HD-cryo SiPM operated at the higher bias voltage of $V_{\mathrm{bias}}$ = 34 V ($\Delta V$ = 6.8 V), clearly showing a shoulder forming in the tail of the 1 PE peak. This artifact is suppressed in the amplitude distribution, since each filled entry in the distribution corresponds to the amplitude from each, individual identified pulse, regardless of whether it originates from DCR or CDAs.
An additional cut is made on the normalised amplitude distribution before determining the mean in order to remove any identified pulses with amplitudes <<1 PE, attributed either to electronic noise or to CDAs that occur on short timescales after the primary avalanche (i.e., before the SiPM has had a chance to recover following a discharge). Using the fitted mean ($\mu_{\mathrm{1PE}}$) and standard deviation ($\sigma_{\mathrm{1PE}}$) from the gaussian fit to the 1 PE peak, we remove any pulses from the distribution with amplitudes $\mu_{\mathrm{1PE}}-3\sigma_{\mathrm{1PE}}$. Figure 9 (left) shows the normalised amplitude distribution at an operating voltage of $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 4.8 V). The purple vertical line indicates the mean of the distribution, corresponding to $\bar{N}_{\mathrm{APA}}$ = 0.170 $\pm$ 0.002. Figure 10 shows $\bar{N}_{\mathrm{APA}}$ as a function of $\Delta V$. As expected, $\bar{N}_{\mathrm{APA}}$ increases as a function of $\Delta V$, increasing by 64% across 4.8 V $\Delta V<$ 6.8 V.
At 77 K, the DiCT probability is reported to be 13% at $\Delta V$ = 5 V for the NUV-HD-cryo technology  [NUV-HD-cryo], whilst our measurement of $\bar{N}_{\mathrm{APA}}$ = 0.170 $\pm$ 0.002 at $\Delta V=$ 4.8 V corresponds to a DiCT probability of 16% $\pm$ 0.2% [^1] The DiCT probability is calculated using the mean number of additional prompt avalanches with the following equation: $P=1\ -\ e^{{}^{-}\bar{N}_{\mathrm{APA}}}$.. Whilst this could suggest a small increase in the DiCT probability when operating at $T_{\mathrm{MC}}$, this difference could also be attributed to small differences between data selection and analysis choices, i.e., our decision to cut on pulse amplitudes of $\mu_{\mathrm{1PE}}-3\sigma_{\mathrm{1PE}}$.
Figure 9: Left: Normalised amplitude distribution at $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 4.8 V), used to determine $\bar{N}_{\mathrm{APA}}$. The purple vertical line indicates the mean of the distribution, corresponding to $\bar{N}_{\mathrm{APA}}$ = 0.170 $\pm$ 0.002. Right: Pulse charge distribution at $V_{\mathrm{bias}}$ = 34 V ($\Delta V$ = 6.8 V), illustrating a clear shoulder forming in the tail in the 1 PE peak due to enhanced CDA probability at higher $\Delta V$.
@@
Figure 10: Mean number of additional prompt avalanches, $\bar{N}_{\mathrm{APA}}$, as a function of $\Delta V$. $\bar{N}_{\mathrm{APA}}$ increases by 64% for 4.8 V $\Delta V<$ 6.8 V. Solid red line is interpolated between the data points.
@@ 3.4 Correlated delayed avalanches
CDAs are a type of correlated noise in SiPMs, comprised of DeCT and afterpulses (AP). DeCT events occur within very short timescales, typically $\mathcal{O}(\mathrm{10\ ns})$ after the primary avalanche, whereas AP events tend to occur on longer timescales, $\mathcal{O}$(ns - µs).
@@
Similarly to DiCT, APs artificially increase the number of detected photons. APs can also distort integration measurements for high photon fluxes, effectively serving as an additional gain factor for the measured signal. The AP probability is expected to increase with $\Delta V$ [NUV-HD-cryo], since a higher $\Delta V$ results in a larger initial avalanche charge. This, combined with an increase in the probability of an avalanche, results in a higher AP probability.
@@
We determine the mean number of CDAs, $\bar{N}_{\mathrm{CDA}}$, within three different time windows following a primary avalanche: 1 µs, 5 µs, and 10 µs. The mean number of CDAs is calculated on a waveform-by-waveform basis. For each waveform, we count how many subsequent pulses are found within a 1 µs, 5 µs, and 10 µs window after the primary pulse (i.e., the first pulse in the waveform that caused the trigger to occur). We apply no requirement on the amplitude of pulses found <<1 µs with respect to the primary pulse, since the gain is decreased for APs occurring within the SiPM recharge time. However, for pulses detected $\geq$1 µs from the primary pulse, we apply a requirement on the pulse amplitude to be $\geq\mu_{\mathrm{1PE}}-3\sigma_{\mathrm{1PE}}$, to reduce leakage from electronic noise. We make one additional requirement, that the time difference between the primary pulse of a waveform $p_{i}$ and the last pulse of the preceding waveform $p_{i-1}$ is $\geq$ 1 s. This cut is made to ensure that the primary pulse originates from DCR or from a photon-induced PE, and is not a CDA of a previous pulse. The choice of 1 s is informed from the unshadowed time distribution in Figure 7, which shows the asymptotic rate from pure DCR at time differences $\geq$1 s.
@@
Figure 11 shows the mean number of CDAs as a function of $\Delta V$, within a 1 µs (red), 5 µs (blue), and 10 µs (green) window following a primary avalanche. The mean number of CDAs is measured to be $\bar{N}_{\mathrm{CDA}}$ = 0.170 $\pm$ 0.002 (1 µs window), 0.420 $\pm$ 0.003 (5 µs window), and 0.560 $\pm$ 0.004 (10 µs window) at $\Delta V$ = 4.8 V. $\bar{N}_{\mathrm{CDA}}$ increases by 102% (1 µs window), 146% (5 µs window), and 172% (10 µs window) between $\Delta V$ = 4.8 V and $\Delta V$ = 6.8 V. The mean number of CDAs increases as a function of $\Delta V$ as expected, driven by the increase in AP probability for higher $\Delta V$. We also observe an increase in the mean number of CDAs for longer time windows following a primary avalanche for a given $\Delta V$ value.
@@
Figure 11: Mean number of CDAs, $\bar{N}_{\mathrm{CDA}}$, within a 1 µs (red), 5 µs (blue), and 10 µs (green) window following a primary avalanche as a function of $\Delta V$. $\bar{N}_{\mathrm{CDA}}$ increases by 102% (1 µs window), 146% (5 µs window), and 172% (10 µs window) between $\Delta V$ = 4.8 V and $\Delta V$ = 6.8 V. Solid red, dashed blue, and dotted green lines are interpolated between the data points. Error bars are shown, however are not necessarily visible as the errors are small.
@@ 3.4.1 Enhanced afterpulsing at low temperatures
At 77 K, the AP probability is reported to be 12% at $\Delta V$ = 5 V [NUV-HD-cryo]. We determine an AP probability of 43% at $\Delta V$ = 4.8 V [^2] The AP probability is calculated using the mean number of correlated delayed avalanches with the following equation: $P=1\ -\ e^{{}^{-}\bar{N}_{\mathrm{CDA}}}$. assuming a 10 µs integration window, as is typically considered for computing AP probabilities in SiPMs [acerbi2015nuv, acerbi2019understanding]. We also assume that the number of DeCT events contributing to the mean number of CDAs we measure is negligible, and that the CDAs we observe originate predominately from AP. We make this assumption based on the timescale of DeCT events, which occur at delay times << 20 ns, which we do not have the ability to resolve from the primary avalanche in our data.
@@
We observe a substantial increase in the AP probability at $T_{\mathrm{MC}}$ compared to the measured value at 77 K. This behaviour has previously been reported at temperatures below 40 K [hanski2025performance, zhang2022scintillation]. It has also been observed that below 40 K, the average time difference between a primary avalanche and AP(s) increases. We also see evidence for an increase in the delay times of APs with respect to the primary avalanche. The $dt$ distribution shown in Figure 7 shows long tails that extend significantly beyond the typical AP timescale of 10 µs. This is indicative of delayed APs that continue to occur even at long timescales, $\mathcal{O}$(ms).
@@
The combination of a higher AP probability and an increase in the average time delay of APs can result in long-lasting AP “trains”, in which APs trigger further APs, causing a self-sustaining cascade of APs that can last for a long time at ultra-low temperatures, up to approximately 1 ms [hanski2025performance]. This effect is driven by the temperature-dependent trapping and release of charge carriers, discussed in [ridley1983lucky]. Figure 12 shows an example waveform trace acquired at $V_{\mathrm{bias}}=$ 34 V ($\Delta V$ = 6.8 V), consistent with a self-sustaining AP train. The presence of such AP trains clearly limits the photon-counting ability of the SiPM, particularly at higher $\Delta V$ values. Whilst delayed APs will also distort integration measurements over $\mathcal{O}$(ms) timescales, Figure 11 shows that the level of distortion caused by delayed APs can be mitigated by using shorter integration windows and operating at a lower $\Delta V$.
@@
Figure 12: Example waveform trace acquired at $V_{\mathrm{bias}}=$ 34 V ($\Delta V$ = 6.8 V), illustrating long-lasting APs consistent with a self-sustaining AP train. Trace shows raw waveform (blue), baseline-subtracted moving-average filtered waveform (orange), pulse finding threshold (dashed red), and identified pulse times (green).
4 Towards a cryogenic cosmic ray muon veto system

### v2

@@ 3.5 Direct crosstalk
Optical crosstalk is a source of correlated noise in a SiPM that occurs due to the emission of a small number of photons ($\sim$10-5 photons/electron) during an avalanche [20], which can trigger avalanches in the surrounding microcells. Optical crosstalk can be split into two categories:
@@
DiCT manifests as a superposition with the 1 PE (single photon) pulse. For example, if a primary avalanche triggers two secondary crosstalk avalanches, the resulting output pulse will be a 3 PE pulse. DiCT can therefore be a limiting factor for applications requiring accurate photon-counting, since it artificially increases the number of detected photons and degrades the signal-to-noise ratio. The DiCT probability is expected to increase with $\Delta V$, since the probability of an avalanche increases with $\Delta V$. In addition, the number of emitted photons is linearly proportional to $\Delta V$, which also contributes to an increase in the DiCT probability. There appears to be a weak dependence on the DiCT probability with temperature for the NUV-HD-cryo SiPM technology, as reported in [17].
To characterise the DiCT of the SiPM, we measure the pulse count rate as a function of threshold within a prompt coincidence window. As the threshold is increased, the count rate decreases in a step-like fashion, leading to a characteristic “staircase” where each step corresponds to excluding events with increasing numbers of fired microcells (1 PE, 2 PE, etc). The rate measured at the top step, $R_{\mathrm{0.5PE}}$, is due to all detectable pulses, whilst the rate measured at the second step, $R_{\mathrm{1.5PE}}$, is due to DiCT pulses. The probability of DiCT, $P_{\mathrm{DiCT}}$, can then be defined as:
$P_{\mathrm{DiCT}}=\frac{R_{\mathrm{1.5PE}}}{R_{\mathrm{0.5PE}}},$ (3.2)
Figure 11 (left) shows the characteristic staircase obtained at $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 6.3 V). By taking the ratio of the blue and red dashed lines, we determine $P_{\mathrm{DiCT}}$ = 15.8%.
For this analysis, we only consider pulses which triggered the waveform, since we know the pulse trigger time. We first plot the distribution of trigger pulse amplitude times to identify the mean time; we only accept pulses for the DiCT analysis with amplitudes that reconstruct within 20 ns of the mean trigger pulse time. This requirement is made to reduce the bias on the DiCT measurement from DeCT pulses occurring very shortly after the primary pulse, in which case the MA filter and pulse finder will merge the pulses. We performed a test in which we: obtained a single 1 PE waveform at $V_{\mathrm{bias}}$ = 32 V; duplicated the waveform about the pulse and superimposed it on the original pulse (at $t_{\mathrm{p}}$ = 0); shifted it in time to study the change in the pulse shape. This is illustrated in Figure 11 (right), which shows that for this particular waveform, two pulses separated by $\leq$ 80 ns will merge and reconstruct with an amplitude surpassing the 1.5 PE threshold.
Figure 12 shows $P_{\mathrm{DiCT}}$ as a function of $\Delta V$. As expected, $P_{\mathrm{DiCT}}$ increases as a function of $\Delta V$, increasing by 49% across 6.3 V $\Delta V<$ 8.3 V. At 77 K, the DiCT probability is reported to be 13% at $\Delta V$ = 5 V for the NUV-HD-cryo technology [17], compared to our measurement of 15.8% at $\Delta V=$ 6.3 V. Our measurement is approximately 20% higher, however we are also operating at $\Delta V$ about 1 V higher; it is plausible that at $\Delta V=$ 5 V, our measurement would be consistent with the value reported in [17]. Thus, based on the single device we have tested, there is no clear temperature dependence on $P_{\mathrm{DiCT}}$ for the NUV-HD-cryo technology.
Figure 11: Left: Pulse count rate as a function of threshold for $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 6.3 V). As the threshold is increased, the count rate decreases in a step-like fashion, leading to a characteristic “staircase”. Right: Study to explore the effect on pulse reconstruction for delayed avalanches closely following a primary, obtained by selecting a 1 PE waveform $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 6.3 V), duplicating and superimposing the waveform onto the original, and shifting it in time to see the time difference at which the reconstructed amplitude drops below the DiCT threshold of 1.5 PE.
@@
Figure 12: $P_{\mathrm{DiCT}}$ as a function of $V_{\mathrm{bias}}$ ($\Delta V$). $P_{\mathrm{DiCT}}$ increases by 49% between 6.3 V $\Delta V<$ 8.3 V. Solid red line is interpolated between the data points. Statistical error bars are shown, however are not necessarily visible as the errors are small.
@@ 3.6 Correlated delayed avalanches
Correlated delayed avalanches (CDAs) are a type of correlated noise in SiPMs, comprised of DeCT and afterpulses (APs). DeCT events occur within very short timescales, typically $\mathcal{O}(\mathrm{10\ ns})$ after the primary avalanche, whereas AP events tend to occur on longer timescales, $\mathcal{O}$(ns - µs).
@@
Similarly to DiCT, APs artificially increase the number of detected photons. APs can also distort integration measurements for high photon fluxes, effectively serving as an additional gain factor for the measured signal. The AP probability is expected to increase with $\Delta V$ [17], since a higher $\Delta V$ results in a larger initial avalanche charge. This, combined with an increase in the probability of an avalanche, results in a higher AP probability.
@@
We determine the mean number of CDAs, $\bar{N}_{\mathrm{CDA}}$, within three different time windows following a primary avalanche: 1 µs, 5 µs, and 10 µs. The mean number of CDAs is calculated on a waveform-by-waveform basis. For each waveform, we count how many subsequent pulses are found within a 1 µs, 5 µs, and 10 µs window after the primary pulse (i.e., the first pulse in the waveform that caused the trigger to occur). We apply no requirement on the amplitude of pulses found <<1 µs with respect to the primary pulse, since the gain is decreased for APs occurring within the SiPM recharge time. This can be seen in Figure 13 (left), which shows the 2D distribution of secondary pulse amplitudes versus $dt$ for $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 6.3 V).
However, we do apply a ToT requirement on pulses detected <<1 µs from the primary pulse to reduce contamination from noise fluctuations occurring on pulse tails. Figure 13 (right) shows the ToT distribution for pulses occurring at $dt\leq$ 500 ns, 500 ns $dt\leq$ 1 µs, and $dt>$ 1 µs with respect to the primary pulse for $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 6.3 V). There is an increase in identified pulses for $dt$ values less than and on the order of the average length of a single photon pulse ($\sim$500 ns), a population which decreases with larger $dt$. We decide to exclude pulses from the CDA analysis with ToT < 80 ns for $dt\leq$ 1 µs; this aligns with the second peak visible in the blue distribution of Figure 13, above which we believe pulses are dominated by real CDAs rather than noise from tail fluctuations.
Figure 13: Left: 2D distribution of secondary pulse amplitudes versus $dt$ for $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 6.3 V), showing recharge-limited APs occurring within $dt<$ 1 µs of a primary pulse. Right: ToT distribution for pulses occurring at $dt\leq$ 500 ns (blue), 500 ns $dt\leq$ 1 µs (orange), and $dt>$ 1 µs (green) with respect to a primary pulse for $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 6.3 V). Red dashed line corresponds to the ToT value above which we accept pulses (with $dt\leq$ 1 µs) for the CDA analysis.
For pulses detected $\geq$1 µs from the primary pulse, we apply a requirement on the pulse amplitude to be $\geq\mu_{\mathrm{SPE,amp}}-3\sigma_{\mathrm{SPE,amp}}$, to reduce leakage from noise. We make one additional requirement, that the time difference between the primary pulse of a waveform $p_{i}$ and the last pulse of the preceding waveform $p_{i-1}$ is $\geq$ 1 s. This cut is made to ensure that the primary pulse originates from DCR or from a photon-induced PE, and is not a CDA of a previous pulse. The choice of 1 s is informed from the unshadowed time distribution in Figure 9, which shows the asymptotic rate from pure DCR at time differences $\geq$1 s.
@@
Figure 14 shows the CDA probability, $P_{\mathrm{CDA}}$, as a function of $\Delta V$ within a 1 µs (red), 5 µs (blue), and 10 µs (green) window following a primary avalanche [^3] The CDA probability is calculated using the mean number of correlated delayed avalanches with the following equation: $P_{\mathrm{CDA}}=1\ -\ e^{{}^{-}\bar{N}_{\mathrm{CDA}}}$.. The CDA probability is measured to be $P_{\mathrm{CDA}}$ = 15.7% $\pm$ 0.2% (1 µs window), 34.0% $\pm$ 0.3% (5 µs window), and 42.3% $\pm$ 0.3% (10 µs window) at $\Delta V$ = 6.3 V. $P_{\mathrm{CDA}}$ increases by 59% (1 µs window), 80% (5 µs window), and 79% (10 µs window) between $\Delta V$ = 6.3 V and $\Delta V$ = 8.3 V. However, we do note that as our pulse finder is unable to resolve pulses that occur (on average) less than approximately 200 ns (Figure 3), there could be a systematic underestimation of $P_{\mathrm{CDA}}$ in our results. $P_{\mathrm{CDA}}$ increases as a function of $\Delta V$ as expected, driven by the increase in AP probability for higher $\Delta V$. We also observe an increase in $P_{\mathrm{CDA}}$ for longer time windows following a primary avalanche for a given $\Delta V$ value.
@@
Figure 14: Correlated delayed avalanche probability ($P_{\mathrm{CDA}}$) as a function of $\Delta V$ considering three time windows following a primary avalanche: 1 µs (red), 5 µs (blue), and 10 µs (green). $P_{\mathrm{CDA}}$ increases by 59% (1 µs window), 80% (5 µs window), and 79% (10 µs window) between $\Delta V$ = 6.3 V and $\Delta V$ = 8.3 V. Solid red, dashed blue, and dotted green lines are interpolated between the data points. Statistical error bars are shown, however are not necessarily visible as the errors are small.
@@ 3.6.1 Enhanced afterpulsing at low temperatures
At 77 K, the AP probability is reported to be 12% at $\Delta V$ = 5 V [17]. We determine a CDA probability of 42% at $\Delta V$ = 6.3 V assuming a 10 µs integration window, as is typically considered for computing AP probabilities in SiPMs [4, 3]. We assume that the number of DeCT events contributing to our $P_{\mathrm{CDA}}$ measurement is negligible; we make this assumption based on the timescale of DeCT events, which occur at $dt<$ 100 ns, which we do not have the ability to resolve from the primary avalanche in our data. Therefore, we assume that our $P_{\mathrm{CDA}}$ measurement is dominated by APs.
@@
We observe a substantial increase in the AP probability at $T_{\mathrm{CSNT}}$ = 9.4 $\pm$ 0.2 mK compared to the measured value at 77 K. This behaviour has previously been reported at temperatures below 40 K [18, 27]. It has also been observed that below 40 K, the average time difference between a primary avalanche and AP(s) increases. We also see evidence for an increase in the delay times of APs with respect to the primary avalanche. The $dt$ distribution shown in Figure 9 shows long tails that extend significantly beyond the typical AP timescale of 10 µs. This is indicative of delayed APs that continue to occur even at much longer timescales.
The combination of a higher AP probability and an increase in the average time delay of APs can result in long-lasting AP “trains”, in which APs trigger further APs, causing a self-sustaining cascade of APs that can last for a long time at ultra-low temperatures, up to approximately 1 ms [18]. This effect is driven by the temperature-dependent trapping and release of charge carriers. During an avalanche, a fraction of carriers can be captured by deep-level traps in the silicon bulk and later released with a characteristic time. As the temperature decreases, the thermal emission rate from these traps is strongly suppressed, leading to much longer de-trapping times. When released, these carriers can retrigger avalanches, producing APs. At sufficiently low temperatures, a single primary avalanche can populate multiple traps with a range of lifetimes, resulting in a sequence of delayed releases and thus a correlated train of APs rather than a single event. This mechanism accounts for both the increased AP probability and the emergence of long AP trains at ultra-low temperature operation.
@@
Figure 15 shows an example waveform trace acquired at $V_{\mathrm{bias}}=$ 34 V ($\Delta V$ = 8.3 V), consistent with a self-sustaining AP train. The presence of such AP trains clearly limits the photon counting ability of the SiPM, particularly at higher $\Delta V$ values. Whilst delayed APs will also distort integration measurements, Figure 14 shows that the level of distortion caused by delayed APs can be mitigated by using shorter integration windows and operating at a lower $\Delta V$.
@@
Figure 15: Example waveform trace acquired at $V_{\mathrm{bias}}=$ 34 V ($\Delta V$ = 8.3 V), illustrating long-lasting APs consistent with a self-sustaining AP train. Trace shows raw waveform (blue), baseline-subtracted moving-average filtered waveform (orange), pulse finding threshold (dashed red), and identified pulse times (green).
4 Towards a cryogenic cosmic-ray muon veto system